OpenStudy (darkbluechocobo):

The distance that an object falls varies directly as the square of the time it has been falling. An object falls 75 feet in 4 seconds. Find how far it will fall in 10 seconds? Approximate your answer to the nearest foot.

OpenStudy (darkbluechocobo):

75=k4^2 k=75/16 (75/16)*10^2 4.6875x100= 468.75 469?